OpenStudy (anonymous):

Solve. 3x2 – 7x – 2 = 0

OpenStudy (pfenn1):

\[3x^2-7x-2=(3x+a)(x+b)\] where (a + 3b)=-7 and ab = -2

OpenStudy (anonymous):

\[ \frac{-b\pm\sqrt{b^2-4ac}}{2a}\implies\frac{7\pm\sqrt{49+24}}{6} \]

OpenStudy (pfenn1):

On second thought....use the way @nbouscal says to do it.